OpenStudy (anonymous):

implicit differentiation: sin((x^2)(y^2))=x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is this sin [ x^2y^2] = x ? or (sin x^2 ) y^2 = x

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

please make me fan

OpenStudy (anonymous):

make you fan?

OpenStudy (anonymous):

jkwon, let me answer it. i got this

OpenStudy (anonymous):

\[\sin(x^2y^2)=x \] \[\cos(x^2y^2)(2xy^2+x^2(2y)(y')=1\] \[2xy^2+x^2(2y)y'=1/[\cos(x^2y^2)]\]

OpenStudy (anonymous):

grrrr

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

oh wow

OpenStudy (anonymous):

how do you use latex?

OpenStudy (anonymous):

Oh I only did half of it

OpenStudy (anonymous):

could you explain chain rule?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

how do you type latex like type integral x dx

OpenStudy (anonymous):

So you treat the problem like an onion. You do the outer layer and then you go to the second layer, then the third. So the outer layer is the cosine, and the inner layer is the x^2y^2

OpenStudy (anonymous):

So then....

OpenStudy (anonymous):

Oops i mean sine is the outer layer

OpenStudy (anonymous):

yeah i gotcha. so you just derive each in parts depending on parenthesis

OpenStudy (anonymous):

pretty much yeah ^^

OpenStudy (anonymous):

remember the inside is the product

OpenStudy (anonymous):

haha, jk i think you have this in the bud

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

well im a math major, so whatever needs help, I like to solve it

OpenStudy (anonymous):

whats a non archimedean ordered field?

OpenStudy (anonymous):

can you ask that in a separate question? you're distracting him from helping me

OpenStudy (anonymous):

so yeah, the product rule for the x^2y^2

OpenStudy (anonymous):

yuppp

OpenStudy (anonymous):

notgoodatmath, dont speak unless told to

OpenStudy (anonymous):

i have 15 fans, i have solved a hundred math problems

OpenStudy (anonymous):

what is Y'?

OpenStudy (anonymous):

I just want help with this problem, and you aren't helping me. Your're doing the opposite actually.

OpenStudy (sandra):

hey cantorset, use the equation editor in the lower left hand side of the reply box

OpenStudy (sandra):

for latex

OpenStudy (anonymous):

oh ok did you see my comment on n^n / n!

OpenStudy (sandra):

oh did not check heh, will do so

OpenStudy (anonymous):

i dont see that as an argument what you did to show it converges